mykahstadler:

Kathleen had a bag of candy-coated chocolates. Of the 60 candies in the bag, 20% were red, 30% were orange, 30% were green, and 20% were blue. Of the blue ones, 10% had a damaged candy coating. Approximately how many of the blue ones had a damaged candy coating

jhonyy9:

so given that of the 60 candies 20% were blue so like a first step you need calcule how many is the 20% of 60 to calcule this you can use the rule of 3 simple and write in this way 60 mean 100% x mean 20% cross multiplie and calcule the x value x = 20*60/100 = ? so after you get the x value you know that 10% of this are damaged so to calcule the 10% of x you can use the rule of 3 simple and write - let be y the damaged candy number x mean 100% y mean 10% cross multiplie and calcule the y value y = x*10/100 = ? damaged candy hope helped easy understandably